4. You just received $225,000 from an insurance settlement. You have decided to set this money aside and invest it for your retirement. Currently, your goal is to retire 25 years from today. How much more will you have in your account on the day you retire if you can earn an average return of 10.5 percent rather than just 8 percent
1
Expert's answer
2020-04-27T19:35:00-0400

"A=P(1+i)^n"

where,

A=final amount

P=initial amount

n=time periods

i=interest rate


assume yearly compounded the rate,


final value for 10.5% rate,

"A_1=225000(1+0.105)^{25}"

final value for 8% rate,

"A_2=225000(1+0.08)^{25}"


"extra\ earning(E)=A1-A2 \\
E=225000*1.105^{25}-225000*1.08^{25}\\
\bold{E=\$1189576.76}"